# Using Dynamic Safety Margins as Control Barrier Functions

### v4

@@ 2.2 Control Barrier Functions
Remark 2.5.
While we stated Definition 2.4 using a unique class $\mathscr{K}$ function $\alpha$ to simplify the presentation, it is equivalent to considering a different $\alpha_{i}\in\mathscr{K}$ for each $h_{i}$. Tuning each $\alpha_{i}$ allows the user to modulate the performance of the closed-loop system while retaining the safety guarantees as long as (13) is satisfied with $\alpha\to\alpha_{i}$.
As noted in the introduction, finding a CBF for arbitrary constraint sets $\mathcal{X}\subset\mathbb{R}^{n}$ is a challenging task. In fact, Lemma 1 states that, when $\mathcal{X},\mathcal{U}$ are compact, a CBF $h:\mathcal{D}\to\mathbb{R}$ satisfying $\mathcal{X}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}$ exists only if $\mathcal{X}$ is control invariant, which is generally not the case. In spite of this, many practitioners insist on imposing $\mathcal{C}=\mathcal{X}$ to construct a “candidate” CBF (i.e. a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}$ that satisfies $\mathcal{X}=\{\boldsymbol{\mathbf{x}}\in\mathcal{D}\mid h(\boldsymbol{\mathbf{x}})\geq 0\}$) and then tuning a class $\mathscr{K}$ function $\alpha$ on a case-by-case basis. Although this approach can yield good performance, it suffers from the possibility that $\mathcal{K}_{\alpha}(\boldsymbol{\mathbf{x}})=\emptyset$ for some $\boldsymbol{\mathbf{x}}\in\mathcal{X}$, which makes the controller $\beta$ ill-defined.
In this paper, we leverage results from the ERG framework to design valid CBFs that satisfy Definition 2 (and more generally Definition 2.4) and, additionally, are guaranteed to satisfy (9).
2.3 Dynamic Safety Margins
@@
$\forall\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l},\quad f\big(\bar{x}(\boldsymbol{\mathbf{v}})\big)+g\big(\bar{x}(\boldsymbol{\mathbf{v}})\big)\bar{u}(\boldsymbol{\mathbf{v}})=0.$ (15)
Moreover, every equilibrium point $\bar{x}(\boldsymbol{\mathbf{v}})$ is stabilizable.
@@
The vector $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$ is hereafter called a “reference” for the system because it identifies a specific equilibrium point. The set of steady-state admissible references is
$\mathcal{V}=\{\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}\mid\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X},~\bar{u}(\boldsymbol{\mathbf{v}})\in\mathcal{U}\}.$ (16)
@@
In essence, Assumption 1 establishes a continuous mapping from references to stabilizable equilibria, and Assumption 2 states that at least one of these equilibria is safe. By Assumption 1, let $\pi:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}^{m}$, locally Lipschitz continuous on $\mathbb{R}^{n}$ and continuous on $\mathbb{R}^{l}$, be a prestabilizing control law such that $\pi\big(\bar{x}(\boldsymbol{\mathbf{v}}),\boldsymbol{\mathbf{v}}\big)=\bar{u}(\boldsymbol{\mathbf{v}})$, and $\bar{x}(\boldsymbol{\mathbf{v}})$ is an asymptotically stable equilibrium point of the prestabilized system
$\dot{\boldsymbol{\mathbf{x}}}=f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\triangleq f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$ (17)
Given a reference $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, the prestabilized dynamics $f_{\pi}$ ensure the existence of an open set $\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathbb{R}^{n}$ that satisfies $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ and is such that
$\boldsymbol{\mathbf{x}}(0)\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\quad\implies\quad\lim_{t\to\infty}\boldsymbol{\mathbf{x}}(t)=\bar{x}(\boldsymbol{\mathbf{v}}).$ (18)
The set $\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is a region of attraction for $\bar{x}(\boldsymbol{\mathbf{v}})$. Note that, for each reference $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, the input constraint set $\mathcal{U}$ induces state constraints on the prestabilized system. This allows us to define the reference-dependent state constraint set
$\mathcal{X}_{\boldsymbol{\mathbf{v}}}=\{\boldsymbol{\mathbf{x}}\in\mathcal{X}\mid\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}\}.$ (19)
By continuity of $\pi$ on $\mathbb{R}^{l}$, it follows that $\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ is closed in $\mathbb{R}^{n}$.
Let us now define the state-reference space $\mathbb{R}^{n}\times\mathbb{R}^{l}$ obtained by concatenating the state and reference vectors $\boldsymbol{\mathbf{x}};~\boldsymbol{\mathbf{v}}]$.
@@
$\tilde{\mathcal{D}}\triangleq\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}\mid\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\},$ (20)
@@
$\tilde{\mathcal{X}}\triangleq\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}\mid\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}},~\boldsymbol{\mathbf{v}}\in\mathcal{V}\},$ (21)
@@
Proof 2.7.
Let $\{(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\}\subset\tilde{\mathcal{X}}$ be a sequence that converges to $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathbb{R}^{n}\times\mathbb{R}^{l}$. Since $\{\boldsymbol{\mathbf{x}}_{n}\}\subset\mathcal{X}$ and $\{\boldsymbol{\mathbf{v}}_{n}\}\subset\mathcal{V}$, and $\mathcal{X}$, $\mathcal{V}$ are closed, it follows that $\boldsymbol{\mathbf{x}}\in\mathcal{X}$ and $\boldsymbol{\mathbf{v}}\in\mathcal{V}$. Define the sequence $\{\pi(\boldsymbol{\mathbf{x}}_{n},\boldsymbol{\mathbf{v}}_{n})\}\subset\mathcal{U}$ and note that by continuity of $\pi$, it converges to $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$ because $\mathcal{U}$ is closed. With this, we have that $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ and it follows that $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{X}}$.
The following DSM definition satisfies the one proposed in [24, Definition 1], although it makes stronger assumptions by requiring a) differentiability rather than continuity and b) invariance rather than strong returnability.
Definition 2.8.
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be continuously differentiable and define the set $\tilde{\mathcal{C}}=\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{D}}\mid\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0\}$. The function $\Delta$ is called a dynamic safety margin (DSM) for $\pi$ if the following conditions hold
@@
$\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0~\implies~\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0,$ (22c)
@@
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ \min_{i\in\{1,\ldots,n_{c}\}}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\end{bmatrix}\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}}),$ (23)
where $\rho:\mathbb{R}^{l}\times\mathbb{R}^{l}\to\mathbb{R}^{l}$ is any locally Lipschitz continuous function such that $\boldsymbol{\mathbf{r}}\in\mathbb{R}^{l}$ is an asymptotically stable equilibrium point of the system $\dot{\boldsymbol{\mathbf{v}}}(t)=\rho\big(\boldsymbol{\mathbf{v}}(t),\boldsymbol{\mathbf{r}}\big)$. In the ERG literature, $\rho$ is called a navigation field [24, Definition 2]. System (23) ensures that, if $\big(\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big)\in\tilde{\mathcal{C}}$, then $\big(\boldsymbol{\mathbf{x}}(t),\boldsymbol{\mathbf{v}}(t)\big)\in\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}}$ at all times [24]. The ERG literature provides systematic tools for computing DSMs for various classes of systems [32, 24, 33, 34, 35, 36]. The following section shows how these tools from the ERG literature can be used to construct valid CBFs.
3 From Dynamic Safety Margins to Control Barrier Functions
@@
3.1 DSMs are CBFs
Based on the intuition that both CBFs and DSMs can be used to design add-on modules for constraint handling, the following theorem states that, a dynamic safety margin is, in fact, a CBF for an augmented system consisting of the concatenation of $\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}$ and $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\boldsymbol{\mathbf{u}}\\ \boldsymbol{\mathbf{w}}\end{bmatrix}$, (24)
where $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathbb{R}^{m}\times\mathbb{R}^{l}$ are the augmented inputs.
Theorem 3.9 (DSMs are CBFs).
If $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ is a DSM for $\pi$, then $\Delta$ is a (vector-valued) CBF with respect to the augmented system (24).
Proof 3.10.
@@ 3.1 DSMs are CBFs
$\tilde{\mathcal{C}}_{i}(c)=\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}:0\leq\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq c\},$ (25)
@@
$\hat{\alpha}_{i}(c)=-\inf_{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(c)}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$ (26)
where $i\in\{1,\ldots,n_{c}\}$. Note that for any $c\in[0,\infty)$, the set $\tilde{\mathcal{C}}_{i}(c)\subset\tilde{\mathcal{C}}$ is compact. It can also be shown that the set-valued maps $\tilde{\mathcal{C}}_{i}$ are both upper- and lower-hemicontinuous. Thus, it follows by Berge’s maximum theorem [31] that each $\hat{\alpha}_{i}$ is continuous on $0,\infty)$. Moreover, it is nondecreasing because for any $c_{1}\leq c_{2}$, $\tilde{\mathcal{C}}_{i}(c_{1})\subset\tilde{\mathcal{C}}_{i}(c_{2})$. Next, note that for any $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(0)$, $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$, and it follows by (22c) that
$\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0.$
@@
$\hat{\alpha}_{i}(0)=-\inf_{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(0)}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq 0.$ (27)
There always exists $\alpha_{i}\in\mathscr{K}$ such that $\forall c\in[0,\infty)$, $\hat{\alpha}_{i}(c)\leq\alpha_{i}(c)$. Define $\alpha(c)=\max_{i\in\{1,\ldots,n_{c}\}}\alpha_{i}(c)$ and note that $\alpha\in\mathscr{K}$. Let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ and $i\in\{1,\ldots,n_{c}\}$ be given and note that since $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$, it follows that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$. With this,
$\dot{\Delta}_{i}\big(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big)$ $=\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})+\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{v}}}0$
$\geq\inf_{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}))}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$
$\geq-\hat{\alpha}_{i}\big(\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big)$
$\geq-\alpha_{i}\big(\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big)$
$\geq-\alpha\big(\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big).$
We have shown that $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})=\big(\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big)\in\mathcal{U}\times\mathbb{R}^{l}$ is such that
$\min_{i\in\{1,\ldots,n_{c}\}}\left[\dot{\Delta}_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})+\alpha\big(\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big)\right]\geq 0.$ (28)
We conclude that the channels $\Delta_{i}$ have the control-sharing property, so $\Delta$ is a (vector-valued) CBF.
Let us define the set of viable inputs at $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ for DSM-CBFs as
@@
The following corollary states that DSM-CBFs satisfy (9), so $\mathcal{U}$ needs not be compact to conclude control invariance of $\tilde{\mathcal{C}}$.
Corollary 3.11.
@@
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$. Then, there exists $\alpha\in\mathscr{K}$ such that $\Delta$ satisfies (9). Specifically, $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\big(\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big)\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. Moreover, if 0 is a regular value of each $\Delta_{i}$, then the set $\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}}\subset\mathcal{X}\times\mathcal{V}$ is control invariant.
Proof 3.12.
The first statement was already shown in the proof of Theorem 3.9. The control invariance of $\tilde{\mathcal{C}}$ is an immediate consequence of Theorem 2 when $n_{c}=1$. For the general case $n_{c}>1$, we refer the reader to [30, Proposition 4.3.7].
@@
A significant concern in the CBF literature arises when the term $L_{g}h(\boldsymbol{\mathbf{x}})=0$ for some $\boldsymbol{\mathbf{x}}\in\mathcal{C}$. This means $h$ lacks uniform relative degree 1 over $\mathcal{C}$, and implies a loss of controllability that can be catastrophic when $L_{f}h(\boldsymbol{\mathbf{x}})<-\alpha\big(h(\boldsymbol{\mathbf{x}})\big)$. Although many relevant works address this issue (e.g. [7, 8]), and Corollary 3.11 already shows it is not a concern for DSM-CBFs, the following corollary shows an additional desirable property of DSM-CBFs under total controllability loss.
Corollary 3.13.

### v5

@@ II-B Control Barrier Functions
As noted in the introduction, finding a CBF for arbitrary constraint sets $\mathcal{X}\subset\mathbb{R}^{n}$ is a challenging task. In fact, Theorem 3 states that a CBF $h:\mathcal{D}\to\mathbb{R}^{n_{c}}$ with superlevel set $\mathcal{X}$ exists only if $\mathcal{X}$ is control invariant, which is generally not the case. In spite of this, many practitioners insist on imposing $\mathcal{C}=\mathcal{X}$ to construct a “candidate” CBF (i.e. a continuously differentiable function $h:\mathcal{D}\to\mathbb{R}^{n_{c}}$ that satisfies $\mathcal{X}=\mathcal{C}$ with $\mathcal{C}$ as in (5)) and then tuning a class $\mathscr{K}$ function $\alpha$ on a case-by-case basis. Although this approach can yield good performance, it suffers from the possibility that $\mathcal{K}_{\alpha}(\bm{x})=\emptyset$ for some $\bm{x}\in\mathcal{X}$, which makes the controller $\beta$ ill-defined.
In this paper, we leverage results from the ERG framework to design valid CBFs that satisfy Definition 2.
II-C Dynamic Safety Margins
@@
$\forall\bm{v}\in\mathbb{R}^{l},\quad f\big(\bar{x}(\bm{v})\big)+g\big(\bar{x}(\bm{v})\big)\bar{u}(\bm{v})=0.$ (12)
Moreover, every equilibrium point $\bar{x}(\bm{v})$ is stabilizable.
@@
The vector $\bm{v}\in\mathbb{R}^{l}$ is hereafter called a “reference” for the system because it identifies a specific equilibrium point. The set of steady-state admissible references is
$\mathcal{V}=\{\bm{v}\in\mathbb{R}^{l}\mid\bar{x}(\bm{v})\in\mathcal{X},~\bar{u}(\bm{v})\in\mathcal{U}\}.$ (13)
@@
In essence, Assumption 1 establishes a continuous mapping from references to stabilizable equilibria, and Assumption 2 states that at least one of these equilibria is safe. By Assumption 1, let $\pi:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}^{m}$, locally Lipschitz continuous on $\mathbb{R}^{n}$ and continuous on $\mathbb{R}^{l}$, be a prestabilizing control law such that $\pi\big(\bar{x}(\bm{v}),\bm{v}\big)=\bar{u}(\bm{v})$, and $\bar{x}(\bm{v})$ is an asymptotically stable equilibrium point of the prestabilized system
$\dot{\bm{x}}=f_{\pi}(\bm{x},\bm{v})\triangleq f(\bm{x})+g(\bm{x})\pi(\bm{x},\bm{v}).$ (14)
Given a reference $\bm{v}\in\mathbb{R}^{l}$, the prestabilized dynamics $f_{\pi}$ ensure the existence of an open set $\mathcal{D}_{\bm{v}}\subset\mathbb{R}^{n}$ that satisfies $\bar{x}(\bm{v})\in\mathcal{D}_{\bm{v}}$ and is such that
$\bm{x}(0)\in\mathcal{D}_{\bm{v}}\quad\implies\quad\lim_{t\to\infty}\bm{x}(t)=\bar{x}(\bm{v}).$ (15)
The set $\mathcal{D}_{\bm{v}}$ is a region of attraction for $\bar{x}(\bm{v})$. Note that, for each reference $\bm{v}\in\mathbb{R}^{l}$, the input constraint set $\mathcal{U}$ induces state constraints on the prestabilized system. This allows us to define the reference-dependent state constraint set
$\mathcal{X}_{\bm{v}}=\{\bm{x}\in\mathcal{X}\mid\pi(\bm{x},\bm{v})\in\mathcal{U}\}.$ (16)
By continuity of $\pi$ on $\mathbb{R}^{l}$, it follows that $\mathcal{X}_{\bm{v}}$ is closed in $\mathbb{R}^{n}$.
Let us now define the state-reference space $\mathbb{R}^{n}\times\mathbb{R}^{l}$ obtained by concatenating the state and reference vectors $\bm{x};~\bm{v}]$.
@@
$\tilde{\mathcal{D}}\triangleq\{(\bm{x},\bm{v})\in\mathbb{R}^{n}\times\mathbb{R}^{l}\mid\bm{x}\in\mathcal{D}_{\bm{v}}\},$ (17)
@@
$\tilde{\mathcal{X}}\triangleq\{(\bm{x},\bm{v})\in\mathbb{R}^{n}\times\mathbb{R}^{l}\mid\bm{x}\in\mathcal{X}_{\bm{v}},~\bm{v}\in\mathcal{V}\},$ (18)
@@
Proof:
Let $\{(\bm{x}_{n},\bm{v}_{n})\}\subset\tilde{\mathcal{X}}$ be a sequence that converges to $(\bm{x},\bm{v})\in\mathbb{R}^{n}\times\mathbb{R}^{l}$. Since $\{\bm{x}_{n}\}\subset\mathcal{X}$ and $\{\bm{v}_{n}\}\subset\mathcal{V}$, and $\mathcal{X}$, $\mathcal{V}$ are closed, it follows that $\bm{x}\in\mathcal{X}$ and $\bm{v}\in\mathcal{V}$. Define the sequence $\{\pi(\bm{x}_{n},\bm{v}_{n})\}\subset\mathcal{U}$ and note that by continuity of $\pi$, it converges to $\pi(\bm{x},\bm{v})\in\mathcal{U}$ because $\mathcal{U}$ is closed. With this, we have that $\bm{x}\in\mathcal{X}_{\bm{v}}$ and it follows that $(\bm{x},\bm{v})\in\tilde{\mathcal{X}}$. ∎
The following DSM definition satisfies the one proposed in [38, Definition 1], although it makes stronger assumptions by requiring a) differentiability rather than continuity and b) invariance rather than strong returnability.
Definition 3
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be continuously differentiable and define the set $\tilde{\mathcal{C}}=\{(\bm{x},\bm{v})\in\tilde{\mathcal{D}}\mid\Delta(\bm{x},\bm{v})\geq 0\}$. The function $\Delta$ is called a dynamic safety margin (DSM) if the following conditions hold
@@
$\Delta_{i}(\bm{x},\bm{v})=0~\implies~\frac{\partial\Delta_{i}}{\partial\bm{x}}f_{\pi}(\bm{x},\bm{v})\geq 0,$ (19c)
@@
$\begin{bmatrix}\dot{\bm{x}}\\ \dot{\bm{v}}\end{bmatrix}=\begin{bmatrix}f(\bm{x})+g(\bm{x})\pi(\bm{x},\bm{v})\\ \min_{i\in\{1,\ldots,n_{c}\}}\Delta_{i}(\bm{x},\bm{v})\end{bmatrix}\rho(\bm{v},\bm{r}),$ (20)
where $\rho:\mathbb{R}^{l}\times\mathbb{R}^{l}\to\mathbb{R}^{l}$ is any locally Lipschitz continuous function such that $\bm{r}\in\mathbb{R}^{l}$ is an asymptotically stable equilibrium point of the system $\dot{\bm{v}}(t)=\rho\big(\bm{v}(t),\bm{r}\big)$. In the ERG literature, $\rho$ is called a navigation field [38, Definition 2]. System (20) ensures that, if $\big(x(0),v(0)\big)\in\tilde{\mathcal{C}}$, then $\big(x(t),v(t)\big)\in\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}}$ at all times [38]. The ERG literature provides systematic tools for computing DSMs for various classes of systems [37, 38, 24, 39, 29, 16]. The following section shows how these tools can be used to construct valid CBFs.
III From Dynamic Safety Margins to Control Barrier Functions
@@
III-A DSMs are CBFs
Based on the intuition that both CBFs and DSMs can be used to design add-on modules for constraint handling, the following theorem states that a dynamic safety margin is, in fact, a CBF for an augmented system consisting of the concatenation of $\bm{x}\in\mathbb{R}^{n}$ and $\bm{v}\in\mathbb{R}^{l}$
$\begin{bmatrix}\dot{\bm{x}}\\ \dot{\bm{v}}\end{bmatrix}=\begin{bmatrix}f(\bm{x})+g(\bm{x})\bm{u}\\ \bm{w}\end{bmatrix}$, (21)
where $(\bm{u},\bm{w})\in\mathbb{R}^{m}\times\mathbb{R}^{l}$ are the augmented inputs.
Theorem 4 (DSMs are CBFs)
If $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ is a DSM, then $\Delta$ is a CBF with respect to the augmented system (21).
Proof:
@@ III-A DSMs are CBFs
$\tilde{\mathcal{C}}_{i}(c)=\{(\bm{x},\bm{v})\in\tilde{\mathcal{C}}\mid 0\leq\Delta_{i}(\bm{x},\bm{v})\leq c\},$ (22)
@@
$\hat{\alpha}_{i}(c)=-\inf_{(\bm{x},\bm{v})\in\tilde{\mathcal{C}}_{i}(c)}\frac{\partial\Delta_{i}}{\partial\bm{x}}f_{\pi}(\bm{x},\bm{v}),$ (23)
where $i\in\{1,\ldots,n_{c}\}$. Note that for any $c\in[0,\infty)$, the set $\tilde{\mathcal{C}}_{i}(c)\subset\tilde{\mathcal{C}}$ is compact. It can also be shown (see the proof of Theorem 3) that the set-valued maps $\tilde{\mathcal{C}}_{i}$ are both upper and lower hemicontinuous at 0. Thus, it follows by Berge’s maximum theorem [8] that each $\hat{\alpha}_{i}$ is continuous at 0. Moreover, it is nondecreasing because for any $c_{1}\leq c_{2}$, $\tilde{\mathcal{C}}_{i}(c_{1})\subset\tilde{\mathcal{C}}_{i}(c_{2})$. Next, note that for any $(\bm{x},\bm{v})\in\tilde{\mathcal{C}}_{i}(0)$, $\Delta_{i}(\bm{x},\bm{v})=0$, and it follows by (19c) that
$\frac{\partial\Delta_{i}}{\partial\bm{x}}f_{\pi}(\bm{x},\bm{v})\geq 0.$
@@
$\hat{\alpha}_{i}(0)=-\inf_{(\bm{x},\bm{v})\in\tilde{\mathcal{C}}_{i}(0)}\frac{\partial\Delta_{i}}{\partial\bm{x}}f_{\pi}(\bm{x},\bm{v})\leq 0.$ (24)
There always exists $\alpha_{i}\in\mathscr{K}$ such that $\forall c\in[0,\infty)$, $\hat{\alpha}_{i}(c)\leq\alpha_{i}(c)$. Define $\alpha(c)=\max_{i\in\{1,\ldots,n_{c}\}}\alpha_{i}(c)$ and note that $\alpha\in\mathscr{K}$. Let $(\bm{x},\bm{v})\in\tilde{\mathcal{C}}$ and $i\in\{1,\ldots,n_{c}\}$ be given and note that since $\bm{x}\in\mathcal{X}_{\bm{v}}$, it follows that $\pi(\bm{x},\bm{v})\in\mathcal{U}$. With this,
$\dot{\Delta}_{i}\big(\bm{x},\bm{v},\pi(\bm{x},\bm{v}),0\big)$ $=\frac{\partial\Delta_{i}}{\partial\bm{x}}f_{\pi}(\bm{x},\bm{v})+\frac{\partial\Delta_{i}}{\partial\bm{v}}0$
$\geq\inf_{(\bm{x},\bm{v})\in\tilde{\mathcal{C}}_{i}(\Delta_{i}(\bm{x},\bm{v}))}\frac{\partial\Delta_{i}}{\partial\bm{x}}f_{\pi}(\bm{x},\bm{v})$
$\geq-\hat{\alpha}_{i}\big(\Delta_{i}(\bm{x},\bm{v})\big)$
$\geq-\alpha_{i}\big(\Delta_{i}(\bm{x},\bm{v})\big)$
$\geq-\alpha\big(\Delta_{i}(\bm{x},\bm{v})\big).$
We have shown that $\forall(\bm{x},\bm{v})\in\tilde{\mathcal{C}}$, $(\bm{u},\bm{w})=\big(\pi(\bm{x},\bm{v}),0\big)\in\mathcal{U}\times\mathbb{R}^{l}$ is such that
$\min_{i\in\{1,\ldots,n_{c}\}}\left[\dot{\Delta}_{i}(\bm{x},\bm{v},\bm{u},\bm{w})+\alpha\big(\Delta_{i}(\bm{x},\bm{v})\big)\right]\geq 0.$ (25)
We conclude that $\Delta$ is a CBF. ∎
Let us define the set of viable inputs at $(\bm{x},\bm{v})$ for DSM-CBFs
@@
The following result follows from Theorem 3 and Theorem 4.
Corollary 1
@@
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM. If $\Delta$ satisfies the MFCQ, then the set $\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}}\subset\mathcal{X}\times\mathcal{V}$ is control invariant.
@@
A significant concern in the CBF literature arises when the term $L_{g}h(\bm{x})=0$ for some $\bm{x}\in\mathcal{C}$. This means $h$ lacks uniform relative degree 1 over $\mathcal{C}$, and implies a loss of controllability that can be catastrophic when $L_{f}h(\bm{x})<-\alpha\big(h(\bm{x})\big)$. Although many relevant works address this issue (e.g. [36, 10]), the following corollary shows an additional desirable property of DSM-CBFs under total controllability loss.
Corollary 2